OpenStudy (anonymous):

Mrs Simms told her class to find two consecutive even integers such that twice the lesser of 2 integers is 4 less than 2 times the greater integer. Write the equation.

OpenStudy (anonymous):

so?

OpenStudy (anonymous):

small=x large=x+2 2(x)+4=2(x+2)